OpenStudy (anonymous):

Solve by factoring: 3x^2+2x-1=0

OpenStudy (anonymous):

3x^2 + 3x - x - 1 = 0 3x(x-1) + 1(x-1) = 0 (3x+1)(x-1) = 0 x = -1/3 ; 1

OpenStudy (anonymous):

Write the solution in set brackets without spaces like: {1,2} how would it now be written?

OpenStudy (anonymous):

?